# M3DHMR: Monocular 3D Hand Mesh Recovery

### v1

@@ 4.3 Ablation Study
As illustrated in Tab. 2, Dynamic Spiral Convolution decreases PJ and PV of the 3D decoder and dramatically increases F@5 compared with SpiralConv++ and Depth-Separable SpiralConv. The results demonstrate that adaptively adjusting the convolution weights for vertices at different positions significantly enhances the reconstruction outcomes.
@@
Tab. 3 demonstrates the effect of DSC Layers at different values of $K$. It can be observed that the model exhibits optimal performance with $K$=8. A value that is too small may result in limited expression and insufficient adaptive ability. Conversely, a value that is too large may lead to an excessive focus on detailed features, which may result in the neglect of some main features and an adverse effect on the model’s generalization ability.

### v2

@@ 4.3 Ablation Study
As illustrated in Tab. 2, Dynamic Spiral Convolution decreases PJ and PV of the 3D decoder and increases F@5 compared with SpiralConv++ and Depth-Separable SpiralConv. The results demonstrate that adaptively adjusting the convolution weights for vertices at different positions significantly enhances the reconstruction outcomes.
@@
Tab. 3 demonstrates the effect of DSC Layers at different values of $K$. It can be observed that the model exhibits optimal performance with KK=8. A value that is too small may result in limited expression and insufficient adaptive ability. Conversely, a value that is too large may lead to an excessive focus on detailed features, which may result in the neglect of some main features and an adverse effect on the model’s generalization ability.